• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name V. Sulkosky

    Log entry time 00:12:00 on April21,2010

    Entry number 318542

    keyword=Swing Shift Summary (better format)

    April 20, 2010
    SL: V. Sulkosky
    TO: M. Canan
    3rd: J. Han Lee
    Swing Shift Summary:
    ++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++

    Plan: Compton tuning and then resume production running.
    16:00 Begin shift waiting for beam recovery after beam studies.
    16:25 MCC called. They are trouble-shooting a vacuum valve in the South LINAC. No time estimate.
    16:55 MCC ready to send tune beam to check out Hall A line.
    17:10 Beginning beam recovery procedure after long down. But MCC is doing some work for Hall B first.
    17:38 Jay Benesch is ready for Compton tuning as soon as beam is restored.

    18:04 MCC called and will attempt to send tune beam into Hall A.
    18:08 Compton tuning begins.
    18:10 ROC 23 crashed. CODA recovery in progress.
    18:40 Noticed that left Q1 current was at 0 Amps and all left arm regulators were off.
    18:50 Jay Benesch called and said they had multiple BLM trips at 1C12 after restoring optics settings from the weekend.
    They are working to fix this problem.
    19:00 Left HRS magnets are ready.
    19:46 MCC sending tune beam to check Hall A line.
    19:56 Talked with Jay. 2L21 was causing some problems and was bypassed.
    Jay will now spend a bit of time working on the Compton tune.
    20:30 Jay called and said either we can have better Compton rates but poor dithering or vice-versa.
    He will consult with the RC to see which is more important.
    20:40 Compton S/N ratio is about 6:1. Dithering status is unclear at this moment.
    20:50 Beginning beam recovery procedure after long down.
    20:54 Ask for Raster on: 17 X 7 in MCC units. Raster is about 4 X 3 mm2.
    21:02 Move to BeO to check visual spot on camera.
    21:07 Target in position. Ask for 10 uA.
    21:11 Ask for 5 nA, Slit to 65 and attenuator to 13. Turn L-HRS upper scintillator HV on. Rate was 2.2 kHz.
    21:15 Move to Lead/Diamond target #3.
    21:17 Target in position. Rate check on Pb/diamond = 12.5 kHz.
    21:24 Turn off upper scintillator HV for left HRS.
    21:27 Ask MCC to ramp current up to 50 uA, 10 uA per minute.
    21:35 Begin production running at 50 uA on lead/diamond target.
    22:36 Stopped Compton run and increased GSO HV to -1650 V. Signal to noise ~ 4.5.
    23:05 Left Q1 current tripped off. Fault was 'max current set'.
    23:15 Hall B needs another 30 minutes before changing IHWP.
    23:30 Ask for IHWP to be retracted.
    23:45 Resume production after DAQ crashes.




    Parity Run Summary:






















     PREX Run #   Beam Current   # of Good Multiplets   Cummulative # of Good Multiplets   Charge (mC)   Cummulative Charge (mC)   Comments   IHWP 
    2979 ---- N/A N/A N/A N/A Comics IN
    2980 ---- N/A N/A N/A N/A Cosmics IN
    2981 ---- N/A N/A N/A N/A Short IN
    2982 ---- N/A N/A N/A N/A Junk IN
    2983 ---- N/A N/A N/A N/A Comics IN
    2984 ---- N/A N/A N/A N/A Junk IN
    2985 ---- N/A N/A N/A N/A Compton tuning (w/ beam) IN
    2986 ---- N/A N/A N/A N/A Comics/beam tuning IN
    2987 ---- N/A N/A N/A N/A Compton tuning (w/ beam) IN
    2988 10-50 uA N/A N/A N/A N/A Beam Ramp IN
    2989 50 uA 88.9 k 1021 k 161.6 1818.5 Production IN
    2990 50 uA 79.9 k 1101 k 145.2 1963.7 Production, left Q1 tripped off IN
    2991 50 uA 13.3 k 1114 k 24.1 1987.8 Production IN
    2992 50 uA 6.3 k 1114 k 11.0 1998.0 Short run IN
    2993 50 uA N/A N/A N/A N/A Junk OUT
    2994 50 uA N/A N/A N/A N/A CODA Crash Junk OUT
    2995 50 uA N/A N/A N/A N/A CODA Crash Junk OUT
    2996 50 uA ---- ---- ---- ---- Production OUT




    HRS Run Summary:







     Left Run #   Beam Current   Comments 
    27451 ---- Junk wrong prescales
    27452 ---- Test run T8 only
    27453 5.0 uA Raster Check T8 only
    27454 5.0 nA Raster Check T5 Lead/diamond





    Compton Run Summary:








     Run #   Start Time   Comments   IHWP 
    22573 21:36 Production (-1.6 kV) IN
    22574 22:36 Production (-1.65 kV) IN
    22575 23:27 Short Test Run IN
    22576 23:38 Production (-1.65 kV) OUT